Find the equations of the tangents and normal to the following curves:


  1. y2 + 8x = 0, parallel to x + y + 4 = 0.
  2. x2 = 3y, perpendicular to x – 2y + 7 = 0.
  3. x2 + 9y2 = 25, parallel to 4x + 9y + 30 = 0.
  4. 25x2 + 4y2 = 100, perpendicular to 8x – 15y + 4 = 0.
  5. x2 – y2 = 15, parallel to 4x – y + 20 = 0.
1
Expert's answer
2021-11-25T09:15:12-0500

1.

"y^2+8x=0"

"2yy'=-8"

"y'=\dfrac{-4}{y}""slope=m_1=y'=\dfrac{-4}{y}"

Tangent is parallel to "x+y+4=0."


"y=-x-4, slope=m_2=-1"

Then "m_1=m_2=>-\dfrac{4}{y}=-1=>y=4"

Find "x"


"(4)^2+8x=0=>x=-2"

Point "(-2, 4)"

Tangent line is


"y-4=-(x-(-2))"

The equation of the tangent line in slope-intercept form


"y=-x+2"



Find the equation of the normal at the point "(-2, 4)"


"slope=m_3=-\dfrac{1}{m_1}=1"

"y-4=(x-(-2))"

The equation of the normal line in slope-intercept form


"y=x+6"

2.

"x^2=3y"

"2x=3y'"

"y'=\dfrac{2x}{3}""slope=m_1=y'=\dfrac{2x}{3}"

Tangent is  perpendicular to "x-2y+7=0."

"y=\dfrac{1}{2}x+\dfrac{7}{2}, slope=m_2=\dfrac{1}{2}"

Then "m_1=-\dfrac{1}{m_2}=>\dfrac{2x}{3}=-2=>x=-3"

Find "y"

"(-3)^2=3y=>y=3"

Point "(-3, 3)"

Tangent line is


"y-3=-2(x-(-3))"

The equation of the tangent line in slope-intercept form


"y=-2x-3"



Find the equation of the normal at the point "(-3, 3)"

"slope=m_3=-\dfrac{1}{m_1}=\dfrac{1}{2}"

"y-3=\dfrac{1}{2}(x-(-3))"

The equation of the normal line in slope-intercept form


"y=\dfrac{1}{2}x+\dfrac{9}{2}"



3.

"x^2+9y^2=25"

"2x+18yy'=0"

"y'=\dfrac{-x}{9y}""slope=m_1=y'=-\dfrac{x}{9y}"

Tangent is parallel to "4x + 9y + 30 = 0."


"y=-\dfrac{4}{9}x-\dfrac{10}{3}, slope=m_2=-\dfrac{4}{9}"

Then "m_1=m_2=>-\dfrac{x}{9y}=-\dfrac{4}{9}=>x=4y"

Find "y"


"(4y)^2+9y^2=25=>y=\pm1"

Point "(-4, -1)," Point "(4, 1)."


Point "(-4, -1)"

Tangent line is


"y-(-1)=-\dfrac{4}{9}(x-(-4))"

The equation of the tangent line in slope-intercept form


"y=-\dfrac{4}{9}x-\dfrac{25}{9}"



Find the equation of the normal at the point "(-4, -1)"


"slope=m_3=-\dfrac{1}{m_1}=\dfrac{9}{4}"

"y-(-1)=\dfrac{9}{4}(x-(-4))"

The equation of the normal line in slope-intercept form


"y=\dfrac{9}{4}x+8"

Point "(4, 1)"

Tangent line is


"y-1=-\dfrac{4}{9}(x-4)"

The equation of the tangent line in slope-intercept form


"y=-\dfrac{4}{9}x+\dfrac{25}{9}"



Find the equation of the normal at the point "(4, 1)"


"slope=m_3=-\dfrac{1}{m_1}=\dfrac{9}{4}"

"y-1=\dfrac{9}{4}(x-4)"

The equation of the normal line in slope-intercept form


"y=\dfrac{9}{4}x-8"

4.

"25x^2+4y^2=100"

"50x+8yy'=0"

"y'=\dfrac{-25x}{4y}""slope=m_1=y'=-\dfrac{25x}{4y}"

Tangent is perpendicular to "8x \u2013 15y + 4 = 0."


"y=\dfrac{8}{15}x+\dfrac{4}{15}, slope=m_2=\dfrac{8}{15}"

Then "m_1=-\dfrac{1}{m_2}=>-\dfrac{25x}{4y}=-\dfrac{15}{8}=>x=\dfrac{3}{10}y"

Find "y"


"25(\dfrac{3}{10}y)^2+4y^2=100=>y=\pm4"

Point "(-\dfrac{6}{5}, -4)," Point "(\dfrac{6}{5}, 4)."


Point "(-\dfrac{6}{5}, -4)"

Tangent line is


"y-(-4)=-\dfrac{15}{8}(x-(-\dfrac{6}{5}))"

The equation of the tangent line in slope-intercept form


"y=-\dfrac{15}{8}x-\dfrac{25}{4}"



Find the equation of the normal at the point "(-\dfrac{6}{5}, -4)"

"slope=m_3=-\dfrac{1}{m_1}=\dfrac{8}{15}"

"y-(-4)=\dfrac{8}{15}(x-(-\dfrac{6}{5}))"

The equation of the normal line in slope-intercept form


"y=\dfrac{8}{15}x-\dfrac{84}{25}"

Point "(\dfrac{6}{5}, 4)"

Tangent line is


"y-4=-\dfrac{15}{8}(x-\dfrac{6}{5})"

The equation of the tangent line in slope-intercept form


"y=-\dfrac{15}{8}x+\dfrac{25}{4}"



Find the equation of the normal at the point "(\dfrac{6}{5}, 4)"


"slope=m_3=-\dfrac{1}{m_1}=\dfrac{8}{15}"

"y-4=\dfrac{8}{15}(x-\dfrac{6}{5})"

The equation of the normal line in slope-intercept form


"y=\dfrac{8}{15}x+\dfrac{84}{25}"

5.

"x^2-y^2=15"

"2x-2yy'=0"

"y'=\dfrac{x}{y}""slope=m_1=y'=\dfrac{x}{y}"

Tangent is parallel to "4x \u2013 y + 20 = 0."


"y=4x+20, slope=m_2=4"

Then "m_1=m_2=>\dfrac{x}{y}=4=>x=4y"

Find "y"


"(4y)^2-y^2=15=>y=\pm1"

Point "(-4, -1)," Point "(4, 1)."


Point "(-4, -1)"

Tangent line is


"y-(-1)=4(x-(-4))"

The equation of the tangent line in slope-intercept form


"y=4x+15"



Find the equation of the normal at the point "(-4, -1)"


"slope=m_3=-\dfrac{1}{m_1}=-\dfrac{1}{4}"

"y-(-1)=-\dfrac{1}{4}(x-(-4))"

The equation of the normal line in slope-intercept form


"y=-\dfrac{1}{4}x-2"

Point "(4, 1)"

Tangent line is


"y-1=4(x-4)"

The equation of the tangent line in slope-intercept form


"y=4x-15"



Find the equation of the normal at the point "(4, 1)"


"slope=m_3=-\dfrac{1}{m_1}=-\dfrac{1}{4}"

"y-1=-\dfrac{1}{4}(x-4)"

The equation of the normal line in slope-intercept form


"y=-\dfrac{1}{4}x+2"